x = 5 + 2x - 3

Solve the value of x
I do believe the answer is -2 but I do not know how to explain each step we take to get the answer. Can someone help?

Answer by mananth(16946)

x = 5 + 2x - 3
add -2x to both sides
x-2x=5+2x-2x-3
-x=5-3
-x=2
divide both sides bvy -1
x=-2